# Efficient Anomaly Detection with Budget Annotation Using Semi-Supervised Residual Transformer

### v2

@@ 3.1 Three levels of supervision
First of all, let us clarify the experiment settings for the three types of supervision as we introduced in LABEL:subsec:supervision.
@@
In the supervised setting (Sup), we randomly draw anomalous images with block-wise annotations from various types of defects to the train set and remove them from the test set. We also involve the original anomaly-free training samples to learn the model in the setting. Note that this setting strictly follows the data splitting principle in (Zhang et al., 2023a) and (Ding et al., 2022). The proposed methods only used the block-wise labels which are obtained by using Equation LABEL:equ:block_label for this supervision condition.
The semi-supervised setting (Semi) is the proposed supervision condition in this work. The same anomalous images are used as training samples while the block labels of them are determined by the bounding-boxes which are defined in Equation LABEL:equ:block_label_semi. Note that in this case, all the blocks are annotated as either unknown or normal.
@@ 3.2 Implementation details
The layer- and layer- feature maps of the WideResnet- model (Zagoruyko and Komodakis, 2016) (pre-trained on ImageNet-1K) are concatenated with a proper scale adaptation and the yielded “hyper-columns” are smoothed via average pooling in the neighborhoods. We set to generate PCFs. For MVTec-AD (Bergmann et al., 2019) and BTAD (Mishra et al., 2021) datasets, we subsample of the PCFs to get the memory banks , while the subsampling ratio for KolektorSDD2 (Božič et al., 2021) is considering the much larger training set of normal images. Our Swin Transformer model consists of blocks with patch csize , the window size is and the number of heads is set to . We train the unsupervised and supervised models from scratch and employ the unsupervised model to initialize the weights of the semi-supervised model as Section LABEL:subsec:mixmatch introduces. The parameters of the focal loss , , and are set to , , and respectively. The sliding window slides over the PCR tensor with a step and the window size is . and are set to and respectively. To speed up training, we sample of sliding windows in normal images, simulated defective images, and true anomalous images at each training iteration. To compare the final prediction map with the ground-truth label map, it is firstly upscaled to the same size as the ground-truth via bilinear interpolation and then smoothed using a Gaussian of kernel with , as it is done in (Roth et al., 2022).
@@ 3.9 Ablation study
In this section, the most influential modules of SemiREST are evaluated in the manner of an ablation study. The involved modules include: the usage of the PCF Section LABEL:subsubsec:pcf, the Bagging prediction of Swin Transformers Section LABEL:subsubsec:bagging, the -NN augmentation of PCRs Section LABEL:subsubsec:knn, the customized MixMatch algorithm Section LABEL:subsec:mixmatch and the random dropout scheme Section LABEL:subsubsec:dropout, respectively. From Table 7 one can observe a consistent increase in the AD performance as more modules are added to the SemiREST model.
In addition, two element-wise distance functions (see Section LABEL:subsubsec:pcf), i.e., the absolute value of the difference (ABS) and the difference square (Square) are also compared in Table 7. According to the table, the square function outperforms the ABS function within unsupervised and semi-supervised settings where no real defective pixels are seen or labeled. The performance gain might be related to the “feature-selection” property of the square operation, which only focuses on the significant components of the residual vector. On the contrary, when the real defective samples are given for generating PCRs, more useful information can be maintained via the conservative ABS function.

### v3

@@ 2.4 MixMatch and Weak Labels Based on Bounding Boxes
3 Our proposed Method
3.1 Problem setting and three supervision conditions
In the conventional setting of an anomaly detection or localization problem, all the training data are anomaly-free. The learned model is then employed to predict the test images with or without anomalies. We depict this supervision condition in the first row of Figure 2. In contrast, some recently proposed methods (Ding et al., 2022; Zhang et al., 2023a; Yao et al., 2023) involve a few anomalous samples to further unleash the discriminative power of more sophisticated models. They term this new setting as “supervised” (as shown in the second row in Figure 2) and thus refer the conventional AD setting as “unsupervised” with a slightly abusive usage of this machine learning term. Note that in the supervised setting, the original anomaly-free samples are also available in training.
In this work, on the other hand, we propose to replace the original pixel-level labels with the block-wise labels (as shown in gray blocks in the second row of Figure 2), and train a Swin transformer model on those “cheap” labels. To further reduce the annotation cost, a novel bounding-boxes label is also designed in this paper. The bounding boxes are defined as the smallest rotated rectangle that can just cover a defective region. Accordingly, the outside region of the bounding box is always defect-free while the labels of the inner pixels are unknown (can be anomalous or not). We term this supervision setting as “semi-supervised” (as shown in the bottom row of Figure 2) and propose a novel algorithm for leveraging the bounding box labels to achieve similar performance to that obtained in the supervised condition.
3.2 Swin Transformer Bagging with Position-constrained Residuals
The overall inference process of our SemiREST algorithm is illustrated in Figure 3. As it can be seen, there are two stages in the whole process, i.e., the generation of position-constrained residual features (the green box) and the bagging process over sliding Swin Transformers (the yellow box).
3.2.1 Position constrained residual features
As mentioned in Section 2.1, we try to exploit the potential usage of the intermediate product of the patch-matching process in PatchCore (Roth et al., 2022). In this work, the feature-matching difference vector, rather than the Euclidean distance between matched features, is used as the network input. Similar usage of the matching differences can also be found in some recently proposed methods (Ding et al., 2022; Zhang et al., 2023a), but they focus on the residual tensors between two holistic feature maps rather than the residual vectors between two feature vectors. Besides, some new evidence shows that positional information of the comparing patches could improve the AD performance (Bae et al., 2022; Gudovskiy et al., 2022) mainly because of the rough alignment of the test images captured in the industrial environment. Accordingly, we introduce the “positional embedding” concept from the literature of Transformers (Dosovitskiy et al., 2021; Liu et al., 2021) to the PatchCore algorithm: the original patch features are aggregated with their positional features encoded in the Transformer way. In this way, the whole patch feature matching process is constrained by the positional information and the ablation study in Section 4.9 verifies the merit of this constraint.
Mathematically, for a input image , one can extract deep features as
where represents a deep network, which is pre-trained on a large but neutral dataset (ImageNet (Russakovsky et al., 2015) for example); denotes the generated deep feature tensor with feature vectors (); stands for the -th deep feature vector extracted from the tensor on the row-column position and .
For a deep feature , we modify it and generate its “Position-Constrained Feature” (PCF) as
where is termed “position code” (PE) in this paper, with its weight parameter and the function denotes the positional embedding process that calculates the -th element of as
where is the token patch size which will be explained in Section 3.2.2.
Note that Equation 3 mainly follows the positional embedding method used in (He et al., 2022). Better position code may be obtained via a careful design of , but that is out of the scope of this work.
One then can build the memory bank following the method proposed in (Roth et al., 2022) as
where is the memory bank containing PCFs (), represents the sophisticated sampling scheme of PatchCore (Roth et al., 2022), denotes the original feature set consisting PCFs, which are extracted from all the normal images for training.
Finally, given a test PCF , the corresponding “Position constrained Residual” (PCR) can be calculated in the PatchCore manner (Roth et al., 2022) as
with the element-wise square function defined as
Compared with the absolute values of the tensor difference employed in Zhang et al. (2023a), the yielded residual vector in Equation 5 is more sparse and this “feature-selection” property empirically benefits the following training process, as it is shown in Section 4.9.
3.2.2 Bagging over sliding Swin Transformers
The Swin Transformer is adopted as the backbone of our deep learning model with the PCRs as its input features, rather than normal images. Follow the conventional ViT (Dosovitskiy et al., 2021; Liu et al., 2021, 2022), the collection of PCRs is firstly partitioned into several “feature patches” over the “Row-Column” plane. Given the pre-defined patch size , the flattened transformer tokens are stored in the token tensor
with the row number , column number and the channel dimension . Note that here every token vector geometrically corresponds to a image-pixel block on the image input of .
As it is illustrated in Figure 3, a number of “sub-tensors” are extracted from in a sliding-window fashion. Let us define the square windows sliding on the Row-Column plane as:
where stands for the coordinate of the window’s top-left corner; is the size of the square window; and . A sub-tensor is sliced from according to as:
The sub-tensor is then fed into the Swin Transformer model to generate the prediction map as
Note that the output of for each token is actually a -D vector, we directly copy the confidences corresponding to the anomaly class to .
In specific, given a pre-defined sampling step and a fixed size , one can generates square sliding windows and the corresponding prediction maps . Now each element of the final prediction map for the holistic tensor can be calculated as
where and are the row and column indexes of the map element; stands for the set of sliding windows overlapping with location ; represents the set cardinality; the subtractions and are the coordinate translation operations.
From the perspective of machine learning, the anomaly score for each token is achieved by aggregating the outputs of Swin Transformers which are fed with different PCRs. In this way, the noise could be averaging out from the final prediction. In this paper, with a gentle generalization to the original definition, we term this sliding-aggregating strategy as Swin Transformer “Bagging” (Bootstrap Aggregating) (Breiman, 1996). The ablation study in Section 4.9 proves the advantage of the Bagging strategy.
3.2.3 Training the Swin Transformer using focal loss
Considering the normal samples usually dominate the original data distribution in most AD datasets, we employ the focal loss (Lin et al., 2017) to lift the importance of the anomaly class. The focal loss used in this work writes
where and stands for the training sample sets (here are transformer tokens) corresponding to defective () and normal () classes respectively, denotes the anomaly confidence (of the current token) predicted by .
3.3 Two Novel Data Augmentation Approaches
Different from the vanilla vision transformers (Dosovitskiy et al., 2021; Liu et al., 2021, 2022), the input to our model is essentially feature residual vectors. Most conventional data augmentation methods (Shorten and Khoshgoftaar, 2019; Yun et al., 2019; Yang et al., 2022) designed for images can not be directly employed in the current scenario. In this work, we propose two novel data augmentation approaches, designed for the PCR input.
3.3.1 Residuals based on -nearest neighbor
Firstly, recall that a PCR is the difference vector between a query feature and its nearest neighbor retrieved in the memory bank , this input feature of the Swin Transformer is unstable as the neighborhood order is sensitive to slight feature perturbations. In this work, we propose to randomly augment the PCRs based on -NN of the query feature, in every forward-backward iteration. Concretely, the PCR is obtained as
where , , are ’s first, second and third neighbors respectively; is a random variable distributes uniformly on the interval ; and are two constants determining the selection probabilities of nearest neighbors; is a small random noise vector and denotes element-wise multiplications. In practice, Equation 12 is used to randomly generate some augmented copies of a given sub-tensor as we illustrate in Algorithm 1 and Figure 4.
In this way, the model can witness a more comprehensive data distribution of PCRs during training and consequently higher performance is achieved, as it is shown in the ablation study Section 4.9.
3.3.2 Random PCR dropout
Inspired by the classic “dropout” learning trick (Srivastava et al., 2014) and the recently proposed MAE autoencoder (He et al., 2022), we design a simple feature augmentation approach termed “random PCR dropout” for realizing higher generalization capacity. In specific, when training, all tokens in the tensor defined in Equation 8 is randomly reset as
where again denotes the random variable sampled from the uniform distribution ; is the constant controlling the frequency of the reset operation. We found this data augmentation also benefits to the final performance in the experiment part of this paper Section 4.9.
3.3.3 Off-the-shelf augmentation methods for generating fake anomalies
Besides the proposed augmentation methods for PCRs, we also follow the off-the-shelf fake/simulated anomaly generation approach proposed in MemSeg (Yang et al., 2023) to ensure an effective learning process of our discriminative model. Readers are recommended to the original work (Yang et al., 2023) for more details.
3.4 Exploiting unlabeled information with ResMixMatch
A long-standing dilemma existing in anomaly detection is whether to involve the real-world defective samples in the training stage. Different choices lead to two main types of AD tasks, i.e., the mainstream “unsupervised” setting (Defard et al., 2021; Roth et al., 2022; Lei et al., 2023; Hou et al., 2021; Bergmann et al., 2022; Liu et al., 2023a) and the minority ”supervised” setting (Ding et al., 2022; Zhang et al., 2023a; Yao et al., 2023; Božič et al., 2021).111Note that here the term “unsupervised” refers to the label absence of the anomaly, normal images are known to be defect-free. As we analyzed in the introduction part, in practice, the key issue is the annotation time rather than the difficulty of obtaining the defective samples.
To strike the balance between the annotation time and model performances, we propose to
label the realistic defects using bounding boxes which require much less annotation labor
compared with the commonly-used pixel-wise labels in the “supervised” setting
(Ding et al., 2022; Zhang et al., 2023a; Yao et al., 2023; Božič et al., 2021). However, as shown in Figure 1, the bounding
boxes, that jointly cover all the anomaly pixels on the image, can only guarantee the
correctness of the negative (normal) label of the outside region of the boxes. The pixel
labels inside the box union are unknown. Fortunately, this semi-supervised situation is
highly well studied in the machine learning literature (Zhu and Goldberg, 2009; Berthelot et al., 2019b, a; Sohn et al., 2020; Wang et al., 2023). In
this work, we borrow the main concept of the MixMatch method (Berthelot et al., 2019b) for
semi-supervised PCR learning. The novel semi-supervised learning algorithm, termed “ResMixMatch”, is summarized in
Algorithm 1. Note that here we assume that the mini-batch only contains one sub-tensor , for the reason of simplicity.
From the algorithm we can see that, the proposed ResMixMatch algorithm only shares the high-level concept of data-mixing with MixMatch (Berthelot et al., 2019b). Besides the replacement of images with residuals, our ResMixMatch differentiates from MixMatch on the following three factors.
We use the two novel augmentation methods i.e., the -NN augmetation defined in Equation 12 and the random PCR dropout defined in Equation 13 for the MixMatch sample generation. The new augmentation methods are carefully desigend to suit PCRs and the Swin Transformer and thus increase the recognition accuracy according to the experiment.
In this work, all the tokens are not treated independently. They are implicitly linked to their original image positions () and image index. When performing the Swin Transformer , the tokens and the corresponding labels from the same image are reorganized into PCR tensors and label maps, as shown in step- and step- in Algorithm 1. In this way, one can sufficiently use the self-attention mechanism of transformers to achieve higher generalization capacity.
In practice, the Swin Transformer model is firstly pre-trained in an unsupervised fashion with true normal samples and synthetic anomalies. Then we fine-tune the model under the semi-supervision determined by the bounding-box labels using the MixMatch scheme.
3.5 The Overall Training Process of SemiREST
3.5.1 The generation of block-wise labels
As shown in Figure 1, we solve the AD task as a block-wise binary classification problem to significantly reduce annotation costs. Recalling that an input image is mapped into the token tensor via Equations 1 and 7, we define a token vector in as a block which covers a pixel region on the original image with defined in Section 3.2.2.
In this paper, we only predict the block labels. Given the pixel binary label map with , the ground-truth label map of the block is estimated as
where denotes the pixel sets corresponding to the image block; and are the two thresholds determining the signs of labels. If the block is labeled as (as shown in Figure 1), no loss gradient will be back-propagated through the corresponding token when training. In this paper, this block labeling strategy is adopted for both the real defects in the supervised setting and the simulated defects in the unsupervised setting.
In the semi-supervised scenario, the pixel labels are estimated by the bounding boxes. In particular, the pixels outside the union of boxes are labeled as (normal) and other pixels are labeled as (unknown). So in this scenario with . Similarly to Equation 14, the block-wise semi-supervision is defined as
where is the threshold for determining whether the label of the block is unknown.
Note that the proposed bounding-box label is partially inspired by the pioneering works (Tabernik et al., 2020; Božič et al., 2021) that employ bounding box to annotate defective parts. However, those bounding boxes consider all the inside pixels as anomalies and thus usually lead to significant label ambiguities. In contrast, SemiREST uses the bounding box to annotate the normal region and leave other pixels as unknown. This labeling strategy is always correct and the unknown region can be successfully exploited by the proposed semi-supervised learning scheme.
3.5.2 Overview of the training process
In summary, the whole training procedure of our SemiREST is depicted in Figure 4. In practice, the full-supervision part (shown in blue) is adopted for the supervised setting as well as the “unsupervised” setting where simulated anomalies are generated. In the semi-supervised setting, both the semi-supervision part (red) and the full-supervision part (blue) are employed to guarantee high performances. Note that this is not against to our original motivation of reducing the annotation effort because one can always use the synthetic defects with no labeling effort.
@@ 4.1 Three levels of supervision
First of all, let us clarify the experiment settings for the three types of supervision as we introduced in 3.1.
@@
In the supervised setting (Sup), we randomly draw anomalous images with block-wise annotations from various types of defects to the train set and remove them from the test set. We also involve the original anomaly-free training samples to learn the model in the setting. Note that this setting strictly follows the data splitting principle in (Zhang et al., 2023a) and (Ding et al., 2022). The proposed methods only used the block-wise labels which are obtained by using Equation 14 for this supervision condition.
The semi-supervised setting (Semi) is the proposed supervision condition in this work. The same anomalous images are used as training samples while the block labels of them are determined by the bounding-boxes which are defined in Equation 15. Note that in this case, all the blocks are annotated as either unknown or normal.
@@ 4.2 Implementation details
The layer- and layer- feature maps of the WideResnet- model (Zagoruyko and Komodakis, 2016) (pre-trained on ImageNet-1K) are concatenated with a proper scale adaptation and the yielded “hyper-columns” are smoothed via average pooling in the neighborhoods. We set to generate PCFs. For MVTec-AD (Bergmann et al., 2019) and BTAD (Mishra et al., 2021) datasets, we subsample of the PCFs to get the memory banks , while the subsampling ratio for KolektorSDD2 (Božič et al., 2021) is considering the much larger training set of normal images. Our Swin Transformer model consists of blocks with patch csize , the window size is and the number of heads is set to . We train the unsupervised and supervised models from scratch and employ the unsupervised model to initialize the weights of the semi-supervised model as Section 3.4 introduces. The parameters of the focal loss , , and are set to , , and respectively. The sliding window slides over the PCR tensor with a step and the window size is . and are set to and respectively. To speed up training, we sample of sliding windows in normal images, simulated defective images, and true anomalous images at each training iteration. To compare the final prediction map with the ground-truth label map, it is firstly upscaled to the same size as the ground-truth via bilinear interpolation and then smoothed using a Gaussian of kernel with , as it is done in (Roth et al., 2022).
@@ 4.9 Ablation study
In this section, the most influential modules of SemiREST are evaluated in the manner of an ablation study. The involved modules include: the usage of the PCF Section 3.2.1, the Bagging prediction of Swin Transformers Section 3.2.2, the -NN augmentation of PCRs Section 3.3.1, the customized MixMatch algorithm Section 3.4 and the random dropout scheme Section 3.3.2, respectively. From Table 7 one can observe a consistent increase in the AD performance as more modules are added to the SemiREST model.
In addition, two element-wise distance functions (see Section 3.2.1), i.e., the absolute value of the difference (ABS) and the difference square (Square) are also compared in Table 7. According to the table, the square function outperforms the ABS function within unsupervised and semi-supervised settings where no real defective pixels are seen or labeled. The performance gain might be related to the “feature-selection” property of the square operation, which only focuses on the significant components of the residual vector. On the contrary, when the real defective samples are given for generating PCRs, more useful information can be maintained via the conservative ABS function.